# Quantum Droplets of Light in Semiconductor Microcavities

### v1

@@ Abstract
Classical liquids arise from the balance between two competing interatomic forces—a short-range repulsion and a long-range induced-dipole-dipole attraction Van der Waals —and may form dense droplets held together by surface tension. Similar configurations also appear in the quantum realm in the form of dilute nanodroplets of liquid Helium Barranco et al. (2016), the prototypical strongly correlated quantum liquid Leggett (2008). Recently, a novel ultradilute self-bound quantum droplet was predicted to exist in ultracold mixtures of bosonic atoms with tunable interactions Petrov (2015) (see also Bulgac (2002)). This arises due to a delicate competition between a net mean-field (MF) attraction—occuring when the interspecies attraction overcomes the intraspecies repulsion—and repulsive Lee-Huang-Yang (LHY) quantum fluctuations Lee et al. (1957); Larsen (1963). Remarkably, shortly after this was proposed, quantum droplets were experimentally realized in homonuclear Cabrera et al. (2018); Semeghini et al. (2018) and heteronuclear D’Errico et al. (2019) binary mixtures, as well as in dipolar condensates Ferrier-Barbut et al. (2016); Chomaz et al. (2016). This has led to intense theoretical and experimental efforts Petrov , for instance on lowering the dimensionality Petrov and Astrakharchik (2016); Cheiney et al. (2018), analyzing rotational Kartashov et al. (2018); Caldara and Ancilotto (2022) and collisional properties Ferioli et al. (2019); Cavicchioli et al. (2025), considering new stabilizing mechanisms Cappellaro et al. (2017); Chiquillo (2025); Mixa et al. (2025), and even different mixtures Wang et al. (2020); Xu et al. (2022). While quantum droplets have proved to be an extremely active research field, the focus has thus far been limited to ultracold atomic gases.
@@
[…]
@@
The non-interacting Hamiltonian $\hat{H}_{0}$ is diagonalized by the transformations $\hat{L}_{\mathbf{k}\sigma}=C_{\mathbf{k}}\hat{c}_{\mathbf{k}\sigma}+X_{\mathbf{k}}\hat{x}_{\mathbf{k}\sigma}$, $\hat{U}_{\mathbf{k}\sigma}=X_{\mathbf{k}}\hat{c}_{\mathbf{k}\sigma}-C_{\mathbf{k}}\hat{x}_{\mathbf{k}\sigma}$, where $\hat{L}_{\mathbf{k}}$ and $\hat{U}_{\mathbf{k}}$ are the lower (LP) and upper (UP) polariton quasiparticles with dispersions $E_{\rm UP/LP}(\mathbf{k})=\frac{1}{2}\left(\epsilon_{\mathbf{k}}^{\rm C}+\epsilon_{\mathbf{k}}^{\rm X}\pm\sqrt{\left(\epsilon_{\mathbf{k}}^{\rm C}-\epsilon_{\mathbf{k}}^{\rm X}\right)^{2}+\Omega_{\rm R}^{2}}\right)$, while $X_{\mathbf{k}}$/$C_{\mathbf{k}}$ are the Hopfield coefficients which quantify the exciton/photon fraction inside each quasiparticle (their explicit expressions can be found in the Supplemental Material (SM) SM ). In the following, we denote the polariton energies at zero momentum as $\epsilon_{\rm{UP}/\rm{LP}}$.
@@
where the collision energy in the scattering is set by the two interacting lower polaritons Bleu et al. (2020). For the repulsive intraspecies interactions $g_{\uparrow\uparrow}=g_{\downarrow\downarrow}\equiv g$, it is sufficient to expand up to second order and use the standard Born approximation Ciuti et al. (1998); Tassone and Yamamoto (1999) for the underlying interactions $v_{\uparrow\uparrow}=v_{\downarrow\downarrow}\equiv v$. We take $g=4\pi/m_{\rm X}$ since this approximately corresponds to the values for both TMDs and QW semiconductors (within a $\sim$10uncertainty de la Fuente Pico et al. (2025)).
[…]
@@
As shown in Fig. 1, the interspecies interaction can be tuned from attractive ($g_{\uparrow\downarrow}<0$) to repulsive ($g_{\uparrow\downarrow}>0$) by varying the photon detuning $\delta_{0}$ across the biexciton Feshbach resonance $\epsilon_{\rm LP}=-\varepsilon_{\rm B}/2$  [^1] Note that the biexciton resonance essentially coincides with the bipolariton resonance due to the smallness of the cavity photon mass Borri et al. (2003).. For concreteness, we use typical experimental parameters for TMD monolayers. While a collapse instability of the polariton condensate is predicted at the MF level Vladimirova et al. (2010) when $\delta g\equiv g+g_{\uparrow\downarrow}<0$ [yellow region in Figs. 1(b-c)], we show here that beyond-MF quantum fluctuations can stabilize the system into a self-bound droplet phase. Note that since the Born approximation represents an upper bound on $g$ Li et al. (2021), in reality we expect the yellow region to extend to larger negative detunings.
@@
Figure 2: (a-d) Thermodynamic potential as a function of exciton density and pairing field for 4 values of the chemical potential $\delta\mu=\mu-\epsilon_{\rm LP}$: 0 (a), 0.10-0.10- 0.10 meV (b), 0.22-0.22- 0.22 meV (c), 0.40-0.40- 0.40 meV (d). The white dot-dashed contour denotes the zero pressure line, i.e., the zeros of $\Omega$, while the green dashed and red solid curves correspond to the conditions $\partial_{\alpha}\Omega=0$ and $\partial_{\Phi}\Omega=0$, respectively. The filled dots in panels (a) and (b) are the stationary points satisfying both conditions. The empty dot in panel (c) indicates the quantum droplet which appears as a first order transition, signaling phase coexistence with the vacuum (white-red point). Thermodynamic potential as a function of pairing field (e) and exciton density (f) for the same values of $\delta\mu$, obtained by evaluating $\Omega$ along the red solid curve of the corresponding panels (a-d). The detuning is fixed at $\delta_{0}=0$, while all the other system parameters are the same as in Fig. 1.
@@
Here, $E_{\pm,\mathbf{k}}^{a/b}$ are the four Bogoliubov branches of the system and we have introduced $A_{\mathbf{k}}=\epsilon_{\mathbf{k}}^{\rm C}-\mu_{\rm MF}$, $B_{\mathbf{k}}=\epsilon_{\mathbf{k}}^{\rm X}-\mu_{\rm MF}+gn_{0}$ and $C_{\pm}=gn_{0}/2\pm\Phi$. The spectrum of excitations is obtained from the diagonalization of the Bogoliubov Hamiltonian, and analytical expressions of the four branches are found for a symmetric mixture. In the limit of small photon mass $m_{\rm C}\ll m_{\rm X}$, the Bogoliubov spectrum simplifies and the momentum integral in Eq. (Bogoliubov theory with bosonic pairing.—) can be computed analytically SM .
@@
[…]
@@
Figure 3: (a) Phase diagram of detuning versus exciton density in the polariton condensate. The blue solid line bounds the region (QD) where quantum droplets can form, and it corresponds to the saturation density of the quantum droplet in phase coexistence with the vacuum. Outside this region, the system is in a superfluid miscible phase (SF). Quantum droplets only occur in the detuning range 19.97-19.97- 19.97 meV $<\delta_{0}<9.6$ meV, where $\delta g\leq 0$ for our parameters. (b) Absorption spectrum, featuring the four Bogoliubov branches $E_{-,\mathbf{k}}^{a/b}$ (bottom panel) and $E_{+,\mathbf{k}}^{a/b}$ (top panel). The LP and UP dispersions are plotted as white dashed lines. The cavity linewidth has been fixed to $\Gamma=0.025$ meV. (c) Droplet density profiles, obtained from a Gross-Pitaevskii (GP)-like approach (see text), for $\delta_{0}=0.0$ (black, solid line), 3.0-3.0- 3.0 (blue dashed) and 6.0-6.0- 6.0 meV (red dash-dotted). System parameters are as in Fig. 1.
@@
Another key observable of the droplet phase is the Bogoliubov spectrum which can be accessed experimentally in polariton systems Pieczarka et al. (2020); Ballarini et al. (2020); Pieczarka et al. (2022); Claude et al. (2023); Frérot et al. (2023). We plot in Fig. 3(b) the excitation absorption spectrum, which is proportional to the photon spectral function $A(\mathbf{k},\omega)=-\frac{1}{\pi}{\rm Im}\left[G_{\text{C}}(\mathbf{k},\omega+i\Gamma)\right]$wik et al. (2016), where the photon Green’s function is defined in SM and $\Gamma$ is the cavity photon linewidth. We observe that the lowest Bogoliubov branch is gapless, i.e., $E^{a}_{-,\mathbf{k}=0}=0$, as a consequence of the $U(1)$ symmetry breaking of the system. The pairing field in our theory leads to a finite energy gap in the next branch $E^{b}_{-,\mathbf{k}}$, which one expects will be removed once further fluctuations are included. The spectrum in Fig. 3(b) also shows two additional branches which lie higher in energy, as they result from the splitting of the UP branch. A clear experimental signature of the polariton quantum droplet phase is the behavior of the speed of sound of the gapless mode, since it remains constant for particle numbers below the saturation density but increases in the miscible superfluid phase.
@@
where the light-matter coupling is accounted for by the effective mass $M_{*}^{-1}=m_{\rm X}^{-1}+\Omega_{R}^{2}/[4m_{\rm C}(\delta_{0}-\mu)^{2}]$, which is linked to the LP effective mass via $M_{*}\simeq X_{0}^{2}M_{\rm LP}$ Carusotto and Ciuti (2013). Here we have eliminated the dependence on the pairing field using the stationary condition (8.iii). Note that the interplay between the inter- and intraspecies interactions is encoded within the density derivative of the thermodynamic potential which thus results in an effective “trapping potential” SM . Figure 3(c) shows how the density profiles have a flat top, and that the saturation density and droplet radius decreases and increases, respectively, as $|\delta g|$ becomes smaller (i.e., considering more negative cavity detuning). Overall, we find droplet radii of order 10$\sim$ 10∼ 10 $\mu$m, which is compatible with ongoing experiments.
@@
In this work, we have shown that exciton-polaritons can host quantum droplets, the ultimate liquid-like behavior of a quantum fluid of light. The advantage of polariton droplets over their cold-atom counterparts is that they do not suffer from three-body recombination, and there is in principle a greater range of particle numbers that can be accessed. Furthermore, the Bogoliubov excitation spectrum is more readily probed in polariton condensates, thus providing a deeper insight into the broken symmetries and associated Goldstone modes of a Bose mixture. The driven-dissipative nature of the polariton system is not expected to change our results qualitatively for long polariton lifetimes Marchetti and Keeling (2014); Sun et al. (2017); Ballarini et al. (2020) and these effects can also be straightforwardly incorporated in our GP analysis, e.g., through a nonresonant pumping scheme Keeling and Berloff (2008).
Our results suggest that quantum droplets of light are achievable for realistic ranges of parameters in current TMD experiments. Key signs of the droplet phase include an ultra-low-threshold for condensation, the absence of a blueshift with varying pump power, an unchanging Bogoliubov spectrum, and a condensate size of $\sim$ 10 $\mu$m. The observation of quantum droplets would provide an unmistakable signature of the elusive quantum nature of exciton-polariton quasiparticles, thus opening promising avenues into the long-standing challenge of quantum polaritonics Gerace et al. (2019); Liew (2023).
@@ Acknowledgements.
MC warmly acknowledges hospitality at the School of Physics and Astronomy of Monash University, where the present work started. OB, JL and MMP acknowledge support from the Australian Research Council (ARC) Centre of Excellence in Future Low-Energy Electronics Technologies (CE170100039), and from ARC Discovery Projects DP240100569 and DP250103746. MMP is also supported through an ARC Future Fellowship FT200100619. FMM acknowledges financial support from the Spanish Ministry of Science, Innovation and Universities through the “Maria de Maetzu” Programme for Units of Excellence in R&D (CEX2023-001316-M) and from the Ministry of Science, Innovation and Universities MCIN/AEI/10.13039/501100011033, FEDER UE, projects No. PID2020-113415RB-C22 (2DEnLight) and No. PID2023-150420NB-C31 (Q).

### v3

@@ Abstract
Classical liquids arise from the balance between two competing interatomic forces—a short-range repulsion and a long-range induced-dipole-dipole attraction Van der Waals —and may form dense droplets held together by surface tension. Similar configurations also appear in the quantum realm in the form of dilute nanodroplets of liquid Helium Barranco et al. (2016), the prototypical strongly correlated quantum liquid Leggett (2008). Recently, a novel ultradilute self-bound quantum droplet was predicted to exist in ultracold mixtures of bosonic atoms with tunable interactions Petrov (2015) (see also Bulgac (2002)). This arises due to a delicate competition between a net mean-field (MF) attraction—occuring when the interspecies attraction overcomes the intraspecies repulsion—and repulsive Lee-Huang-Yang (LHY) quantum fluctuations Lee et al. (1957); Larsen (1963). Remarkably, shortly after this was proposed, quantum droplets were experimentally realized in homonuclear Cabrera et al. (2018); Semeghini et al. (2018) and heteronuclear D’Errico et al. (2019) binary mixtures, as well as in dipolar condensates Ferrier-Barbut et al. (2016); Chomaz et al. (2016). This has led to intense theoretical and experimental efforts Petrov , for instance on lowering the dimensionality Petrov and Astrakharchik (2016); Cheiney et al. (2018); Lavoine and Bourdel (2021), analyzing rotational Kartashov et al. (2018); Caldara and Ancilotto (2022) and collisional properties Ferioli et al. (2019); Cavicchioli et al. (2025), considering new stabilizing mechanisms Cappellaro et al. (2017); Lavoine et al. (2021); Chiquillo (2025); Mixa et al. (2025), and even different mixtures Wang et al. (2020); Xu et al. (2022). While quantum droplets have proved to be an extremely active research field, the focus has thus far been limited to ultracold atomic gases.
@@
Figure 1: (a) Schematic illustration of the energy scales for one (bottom) and two (top) particles. The hybridization of an exciton (X) and a photon (C) yields upper (U) and lower (L) polariton quasiparticles. Two $\uparrow$ and $\downarrow$ lower polaritons repel (attract) each other when their energy $2\epsilon_{\rm LP}$ is above (below) the energy $\varepsilon_{\rm B}$ of the biexciton ($\uparrow-X\downarrow$). (b) Two-lower-polariton (green) and two-photon (black) energies as a function of detuning $\delta_{0}$. The blue point denotes the Feshbach resonance that occurs when $2\epsilon_{\rm LP}=-\varepsilon_{\rm B}$. (c) Detuning dependence of the intra- (blue) and inter-species (red) exciton-exciton interaction coefficients. The magenta (orange) square is an example of repulsive (attractive) interspecies interactions. The black curve shows $\delta g=g+g_{\uparrow\downarrow}$, and the yellow region denotes the necessary condition $\delta g<0$ for the appearance of quantum droplets. We use parameters inspired by recent experiments with MoSe2 monolayers: $\Omega_{\rm R}=28$ meV, $\varepsilon_{\rm X}=470$ meV Dufferwiel et al. (2015), $\varepsilon_{\rm B}=20$ meV Hao et al. (2017), and exciton mass $m_{\rm X}=1.14m_{0}$ Kylänpää and Komsa (2015) (with $m_{0}$ the bare electron mass).
@@
The non-interacting Hamiltonian $\hat{H}_{0}$ is diagonalized by the transformations $\hat{L}_{\mathbf{k}\sigma}=C_{\mathbf{k}}\hat{c}_{\mathbf{k}\sigma}+X_{\mathbf{k}}\hat{x}_{\mathbf{k}\sigma}$, $\hat{U}_{\mathbf{k}\sigma}=X_{\mathbf{k}}\hat{c}_{\mathbf{k}\sigma}-C_{\mathbf{k}}\hat{x}_{\mathbf{k}\sigma}$, where $\hat{L}_{\mathbf{k}\sigma}$ and $\hat{U}_{\mathbf{k}\sigma}$ are the lower (LP) and upper (UP) polariton quasiparticles with dispersions $E_{\rm UP/LP}(\mathbf{k})=\frac{1}{2}\left(\epsilon_{\mathbf{k}}^{\rm C}+\epsilon_{\mathbf{k}}^{\rm X}\pm\sqrt{\left(\epsilon_{\mathbf{k}}^{\rm C}-\epsilon_{\mathbf{k}}^{\rm X}\right)^{2}+\Omega_{\rm R}^{2}}\right)$, while $X_{\mathbf{k}}$$C_{\mathbf{k}}$ are the Hopfield coefficients which quantify the exciton/photon fraction inside each quasiparticle (their explicit expressions can be found in the Supplemental Material (SM) SM ). In the following, we denote the polariton energies at zero momentum as $\epsilon_{\rm{UP}/\rm{LP}}$.
@@
where the collision energy in the scattering is set by the two interacting lower polaritons Bleu et al. (2020). For the repulsive intraspecies interactions $g_{\uparrow\uparrow}=g_{\downarrow\downarrow}\equiv g$, it is sufficient to expand up to second order and use the standard Born approximation Ciuti et al. (1998); Tassone and Yamamoto (1999) for the underlying interactions $v_{\uparrow\uparrow}=v_{\downarrow\downarrow}\equiv v$. We take $g=4\pi/m_{\rm X}$ since this approximately corresponds to the values for both TMDs and QW semiconductors (within a $\sim$10% uncertainty de la Fuente Pico et al. (2025)).
On the other hand, the underlying interspecies interactions are attractive, i.e., $v_{\uparrow\downarrow}=v_{\downarrow\uparrow}<0$, and they lead to a $\uparrow\downarrow$ biexciton bound state which requires the infinite Born series in Eq. (2). Importantly, in the regime $\Omega_{\rm R}\ll\varepsilon_{\rm X}$ the effective interaction $g_{\uparrow\downarrow}$ takes a universal low-energy form in the vicinity of the biexciton bound state with binding energy $\varepsilon_{\rm B}$ Bleu et al. (2020):
@@
In contrast to 2D ultracold gases, $g_{\uparrow\downarrow}$ depends on the light-matter coupling instead of the system density.
As shown in Fig. 1, the interspecies interaction can be tuned from attractive ($g_{\uparrow\downarrow}<0$) to repulsive ($g_{\uparrow\downarrow}>0$) by varying the photon detuning $\delta_{0}$ across the biexciton Feshbach resonance $\epsilon_{\rm LP}=-\varepsilon_{\rm B}/2$  [^1] Note that the biexciton resonance essentially coincides with the bipolariton resonance due to the smallness of the cavity photon mass Borri et al. (2003).. For concreteness, we use typical experimental parameters for TMD monolayers. While a collapse instability of the polariton condensate is predicted at the MF level Vladimirova et al. (2010) when $\delta g\equiv g+g_{\uparrow\downarrow}<0$ [yellow region in Figs. 1(b-c)], we show here that beyond-MF quantum fluctuations can stabilize the system into a self-bound droplet phase. Note that since the Born approximation represents an upper bound on $g$ Li et al. (2021), in reality we expect the yellow region to extend to larger negative detunings. The precise value of the lower critical detuning can also depend on the finite range of the interactions.
@@
Figure 2: (a-d) Thermodynamic potential as a function of exciton density and pairing field for 4 values of the chemical potential $\delta\mu=\mu-\epsilon_{\rm LP}$: 0 (a), 0.10-0.10 meV (b), 0.22-0.22 meV (c), 0.40-0.40 meV (d). The white dot-dashed contour denotes the zero pressure line, i.e., the zeros of $\Omega$, while the green dashed and red solid curves correspond to the conditions $\partial_{\alpha}\Omega=0$ and $\partial_{\Phi}\Omega=0$, respectively. The filled dots in panels (a) and (b) are the stationary points satisfying both conditions. The empty dot in panel (c) indicates the quantum droplet which appears as a first order transition, signaling phase coexistence with the vacuum (white-red point). Thermodynamic potential as a function of pairing field (e) and exciton density (f) for the same values of $\delta\mu$, obtained by evaluating $\Omega$ along the red solid curve of the corresponding panels (a-d). The detuning is fixed at $\delta_{0}=0$, while all the other system parameters are the same as in Fig. 1.
@@
Here, $E_{\pm,\mathbf{k}}^{a/b}$ are the four Bogoliubov branches of the system and we have introduced $A_{\mathbf{k}}=\epsilon_{\mathbf{k}}^{\rm C}-\mu_{\rm MF}$, $B_{\mathbf{k}}=\epsilon_{\mathbf{k}}^{\rm X}-\mu_{\rm MF}+gn_{0}$ and $C_{\pm}=gn_{0}/2\pm\Phi$. The spectrum of excitations is obtained from the diagonalization of the Bogoliubov Hamiltonian, and analytical expressions of the four branches are found for a symmetric mixture. In the limit of small photon mass $m_{\rm C}\ll m_{\rm X}$, the Bogoliubov spectrum simplifies and the momentum integral in Eq. (Bogoliubov theory with bosonic pairing.—) can be computed analytically SM . Notice that the validity of the Bogoliubov theory improves by moving further from the biexciton resonance, i.e., for negative values of detuning (cf. Fig. 1).
@@
$\text{(i)}\,\Omega|_{\text{QD}}$ $=0,$ $\text{(ii)}\,\partial_{\alpha}\Omega|_{\text{QD}}$ $=0,$ $\text{(iii)}\,\partial_{\Phi}\Omega|_{\text{QD}}$ $=0,$ (8)
@@
Figure 3: (a) Phase diagram of detuning versus exciton density in the polariton condensate. The blue solid line bounds the region (QD) where quantum droplets can form, and it corresponds to the saturation density of the quantum droplet in phase coexistence with the vacuum. Outside this region, the system is in a superfluid miscible phase (SF). Quantum droplets only occur in the detuning range 19.97-19.97 meV $\delta_{0}<9.6$ meV, where $\delta g\leq 0$ for our parameters. (b) Absorption spectrum, featuring the four Bogoliubov branches $E_{-,\mathbf{k}}^{a/b}$ (bottom panel) and $E_{+,\mathbf{k}}^{a/b}$ (top panel), which includes a linear gapless mode. The LP and UP dispersions are plotted as white dashed lines. The cavity linewidth has been fixed to $\Gamma=0.025$ meV. (c) Droplet density profiles, obtained from a Gross-Pitaevskii (GP)-like approach (see text), for $\delta_{0}=0.0$ (black, solid line), 3.0-3.0 (blue dashed) and 6.0-6.0 meV (red dash-dotted). System parameters are as in Fig. 1.
@@
Another key observable of the droplet phase is the excitation spectrum which can be accessed experimentally in polariton systems Pieczarka et al. (2020); Ballarini et al. (2020); Pieczarka et al. (2022); Claude et al. (2023); Frérot et al. (2023). We plot in Fig. 3(b) the excitation absorption spectrum, which is proportional to the photon spectral function $A(\mathbf{k},\omega)=-\frac{1}{\pi}{\rm Im}\left[G_{\text{C}}(\mathbf{k},\omega+i\Gamma)\right]$ Ćwik et al. (2016), where the photon Green’s function is defined in SM and $\Gamma$ is the cavity photon linewidth. We observe that the lowest Bogoliubov branch is linear and gapless, i.e., $E^{a}_{-,\mathbf{k}=0}=0$, as a consequence of the $U(1)$ symmetry breaking of the system. The pairing field in our theory leads to a finite energy gap in the next branch $E^{b}_{-,\mathbf{k}}$, which one expects will be removed once further fluctuations are included. The spectrum in Fig. 3(b) also shows two additional branches which lie higher in energy, as they result from the splitting of the UP branch. A clear experimental signature of the polariton quantum droplet phase is the behavior of the speed of sound of the gapless mode, since it remains constant for particle numbers below the saturation density but increases in the miscible superfluid phase.
@@
[…]
@@
In this work, we have shown that exciton-polaritons can host quantum droplets, the ultimate liquid-like behavior of a quantum fluid of light. Our findings complement previous studies of liquid light in the context of nonlinear optics with competing nonlinearities Michinel et al. (2002, 2006); Novoa et al. (2009); Wu et al. (2013), as well as quasi-equilibrium photon fluids in plasma and semiconductor environments Figueiredo et al. (2024). The advantage of polariton droplets over their cold-atom counterparts is that they do not suffer from three-body recombination, due to the lack of deep two-body bound states Braaten and Hammer (2006), and there is in principle a greater range of particle numbers that can be accessed. Furthermore, the Bogoliubov excitation spectrum is more readily probed in polariton condensates, thus providing a deeper insight into the broken symmetries and associated Goldstone modes of a Bose mixture. Specifically, there is currently a lack of consensus between different perturbative methods Petrov (2015); Hu and Liu (2020) about the dispersions of the spin and density modes. The driven-dissipative nature of the polariton system is not expected to change our results qualitatively for long polariton lifetimes Marchetti and Keeling (2014); Sun et al. (2017); Ballarini et al. (2020) and these effects can also be straightforwardly incorporated in our GP analysis, e.g., through a nonresonant pumping scheme Keeling and Berloff (2008).
[…]
@@ Acknowledgements.
We gratefully acknowledge fruitful discussions with Dmitry Petrov, and we thank Cesar Cabrera for helpful feedback on the manuscript. We also thank Humberto Michinel for making us aware of Refs. Michinel et al. (2002, 2006); Novoa et al. (2009); Wu et al. (2013). MC warmly acknowledges hospitality at the School of Physics and Astronomy of Monash University, where the present work started. OB, JL and MMP acknowledge support from the Australian Research Council (ARC) Centre of Excellence in Future Low-Energy Electronics Technologies (CE170100039), and from ARC Discovery Projects DP240100569 and DP250103746. OB also acknowledges support from the Deutsche Forschungsgemeinschaft (DFG) via the Collaborative Research Centre SFB 1225 ISOQUANT (Project-ID No. 273811115). MMP is also supported through an ARC Future Fellowship FT200100619. FMM acknowledges financial support from the Spanish Ministry of Science, Innovation and Universities through the “Maria de Maetzu” Programme for Units of Excellence in R&D (CEX2023-001316-M) and from the Ministry of Science, Innovation and Universities MCIN/AEI/10.13039/501100011033, FEDER UE, projects No. PID2020-113415RB-C22 (2DEnLight) and No. PID2023-150420NB-C31 (Q). The data that support the findings of this article are openly available Caldara et al. .